OpenStudy (anonymous):

if x^2-y^2=100 and x-y=20, what is the value of x+y?

OpenStudy (dumbcow):

use difference of squares

OpenStudy (anonymous):

\[x^2-y^2 = (x+y)(x-y)= 100 \Rightarrow 20(x+y) = 100\Rightarrow x+y= 5\]